OpenStudy (anonymous):

Can someone check my answers? Find the domain and range of ƒ−1 given f(x) = 3sqrt(x+8+1) Domain: (−9, ∞) and Range: (−1, ∞) Domain: (-∞, ∞) and Range: (-∞, ∞)** Domain: (−8, ∞) and Range: (1, ∞) Domain: (−7, ∞) and Range: (0, ∞)

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hey

OpenStudy (misty1212):

do this: find the domain of \(f\) that will be the range of \(f^{-1}\) then find the range of \(f\) that will be the domain if \(f^{-1}\)

OpenStudy (misty1212):

if you do not know how to find those, i can show you

OpenStudy (er.mohd.amir):

for Domain (x+8+1) greeter and equal to zero

OpenStudy (misty1212):

i am going to guess that there is a typo in the question

OpenStudy (anonymous):

got Domain: (-∞, ∞) and Range: (-∞, ∞). So am I wrong?

OpenStudy (misty1212):

is it \[f(x)=3\sqrt{x+8}+1\] or maybe \[f(x)=\sqrt[3]{x+8}+1\] i doubt it is what you wrote

OpenStudy (anonymous):

\[\sqrt[3]{x+8+1}\]

OpenStudy (anonymous):

The 2nd one. @misty1212

OpenStudy (misty1212):

cubed root right?

OpenStudy (misty1212):

you can take the cubed root of any number so yeah \((-\infty,\infty)\)

OpenStudy (anonymous):

So i'm right?

OpenStudy (misty1212):

yes if it is cubed root

OpenStudy (misty1212):

no one writes \(8+1\) though so i am confused on planet earth it is called 9

OpenStudy (anonymous):

that's how the question is

OpenStudy (misty1212):

ok i believe you

OpenStudy (anonymous):

lol okay

OpenStudy (anonymous):

I don't see why u shouldn't